THE DIMENSION OF THE SPACE OF CUSP FORMS OF WEIGHT ONE

William Duke

Abstract. A new upper bound is given for the dimension of the space of holomorphic cusp forms of weight one and prime level q:

dim S_{1}(q) << q^{11/12} log^{4} q

with an absolute implied constant.

Typeset by AMS-TEX